OpenStudy (anonymous):

PLEASE HELP ME The graph of a line has a slope of 2 and passes through the point (–6, –5). What is the equation of the line in point-slope form? (Points : 4) y – 6 = 2(x – 5) y + 6 = 2(x + 5) y – 5 = 2(x – 6) y + 5 = 2(x + 6) Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. (Points : 4) (y + 1) = 3(x – 2) (y – 1) = 3(x + 2) (x + 2) = 3(y – 1) (x – 2) = 3(y + 1) Write 6x - 3y = 3 in slope-intercept form. Show all work!

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

you can use this equation: y-y1=m(x-x1) where m being the gradient is 2, y1 is -5 and x1 = -6 therefore the answer for the first one is : y-(-5)=2(x-(-6)) therefore: y+5=2(x+6)

OpenStudy (anonymous):

You want to try the second one or i do it for you :) ?

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

sure :) ... for the second one : y-y1=m(x-x1) therefore : m=3 (gradient) , y1=1 and x1=-2 therefor y-1=3(x-(-2) answer= y-1=3(x+2)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

not a problem :) ... can i ask u for some help again pls :) ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

cause i couldnt quite understand the second line of the answer which you gave me :/

OpenStudy (anonymous):

i"m not sure my self i just ask some about your question and they gave me that sorry i wish i could help

OpenStudy (anonymous):

ok :) not a prob :) ... thx for ur help :)

OpenStudy (anonymous):

no problem i wish i could help